The brakes of an automobile can accelerate it at -7.20 m/s2.  What time is required to reduce the velocity of the automobile from 50.0 m/s to 12.5 m/s?



1
Expert's answer
2020-12-14T12:15:09-0500

The time is change in velocity divided by acceleration;


"t=\frac{v_f-v_i}{a}=\frac{12.5-50}{-7.2}=5.2\text{ s}."